jhonyy9 (jhonyy9):

how may be this so difficile to solve it - i like,need asking this,bc. much my college was sure that x=4 is the right answer,... inclusive colleges engineer ,till i got the idea how can i proving that x=4 not is the right value of x - there is the exponential equation 2^16 ------- = 2^2^x 16^2 x = ?

jhonyy9 (jhonyy9):

all colleges was secure,sure that x=4 is the right value of x till i ve wrote this : ok. i said we accept that 4 is right so than there are 2^16 -------- = 2^2^4 16^2 2^16 ------------ = 2^2^4 (2^4)^2 2^16 --------- = 2^2^4 2^8 2^(16-8) = 2^2^4 2^8 = 2^2^4 make logarithm base 2 on both sides log(2) 2^8 = log(2) 2^2^4 using this properties of logarithms log a^x = xlog a so we get 8log(2) 2 = (2^4)log(2) 2 => 8 = 2^4 so 8 = 16 - in this way with this proof was accepted that x=4 not is the right,correct value,answer and and hence was proven that the right correct answer is sure x = 3 bc. 8 = 2^3 q.e.d.

jhonyy9 (jhonyy9):

hi @mathmate wait your opinion about above wrote - and thank you

jhonyy9 (jhonyy9):

i think just a parentheses may be the key of this exercise so than on the left part may be wrote ----------- = (2^2)^x how you think this @mathmate ?

OpenStudy (irishboy123):

\(\dfrac{2^{16}}{16^{2}} = \dfrac{2^{16}}{(2^4)^{2}} = \dfrac{2^{16}}{2^8} = 2^8\) \(2^8 = 2^{2^x} \implies 2^x = 8 \implies x = 3\)

OpenStudy (mathmate):

Well, the question boils down to answering the question What is \(4^{3^2}\) If you evaluated it as \(4^{3^2}=4^9=262144\), then it is correct because exponentiation is right associated. So similarly, 4^3^2 should be evaluated similarly, but noting that 4^3^2 is a sloppy way of writing \(4^{3^2}\) even though they are mathematically the same. Based on the above logic, x=3, as you had it. @jhonyy9 congrats.

OpenStudy (mathmate):

@jhonyy9 2^2^2 should have been written 2^(2^2) for clarity. (2^2)^2 is a different problem and has different results. The rule for chained exponents is to evaluate from right to left, different from +,-,and *.

jhonyy9 (jhonyy9):

i think that my colleges ve tought it in tis way that there are (2^2)^x using the exponential properties that (x^a)^b = x^(a*b) and in this way have got they (2^2)^4 = 2^(2*4) = 2^8 so i just supposed this for i can understanding how they ve got for x this value of 4 like a right answer

OpenStudy (loser66):

\((a^2)^3\neq a^{2^3}\)

OpenStudy (loser66):

I am with @IrishBoy123

jhonyy9 (jhonyy9):

@ganeshie8 @TheSmartOne @Kainui

OpenStudy (kainui):

When you read: a^b^c it means: a^(b^c)

OpenStudy (mathstudent55):